OpenStudy (anonymous):

At Computer Central, the cost of renting a computer is $45 plus $35 per day. The total cost can be represented by the equation y=45+35x, where x is the number of days for which the computer has been rented and y is the total cost of renting the computer. What is the y-intercept of this equation?

OpenStudy (anonymous):

Will give medal!!

OpenStudy (theprofessor35):

The Y intercept is where the line crosses the y axis

OpenStudy (theprofessor35):

Exactly!

OpenStudy (theprofessor35):

y=45+35x y = 45 + 35(0) y = 45 The y intercept is 45

OpenStudy (anonymous):

how do i give medals?

OpenStudy (anonymous):

sorry just started

OpenStudy (theprofessor35):

press best repsonse

OpenStudy (theprofessor35):

on my answer if you'd like :D

OpenStudy (anonymous):

and that gives a medal?

OpenStudy (theprofessor35):

yep.

OpenStudy (anonymous):

how do I become a fan? the same way?

OpenStudy (anonymous):

and thank you so much!

OpenStudy (theprofessor35):

Go to my profile to fan! :)

OpenStudy (anonymous):

ok thanks